OpenStudy (anonymous):

An object is thrown upwards and after 2 seconds is 78 m above where it was initially thrown. Assume no air resistance, and determine the initial speed the object was thrown at.

OpenStudy (istim):

Wrong group.

OpenStudy (amistre64):

right group ... kinda

OpenStudy (istim):

Sorta. Physics.

OpenStudy (amistre64):

we are in meters so -4.9 assume 0 for initial height

OpenStudy (rogue):

\[s(t) = - \frac {1}{2} g t^2 + v_0 t + s_0\]

OpenStudy (amistre64):

-4.6 t^2 + V t = 78

OpenStudy (amistre64):

6, 9, not much difference eh

OpenStudy (istim):

Just use the Kinematics equations, right?

OpenStudy (anonymous):

what does 78m stand for? displacement, distance, or velocity?

OpenStudy (amistre64):

kinetics are for babies and old ladies ....

OpenStudy (istim):

Kinematics aren't kinetics. But okay.

OpenStudy (istim):

78m has no direction, and also speed is m/s.

OpenStudy (anonymous):

but it says upwards :s

OpenStudy (istim):

If it says upwards, then its displacement.

OpenStudy (amistre64):

 it might be easier to assume its a vertex

OpenStudy (amistre64):

a minimum velocity perhaps :)

OpenStudy (anonymous):

kk. thanks what is -4.9? how did we get that?

OpenStudy (amistre64):

-4.9(4) + V(2) - 78 = 0 V = 39 +4.9(2) maybe -4.9 is gravity holding us back

OpenStudy (amistre64):

since we are in meters

OpenStudy (istim):

Wait. I'm stupid. Isn't gravity 9.8?

OpenStudy (anonymous):

yeah gravity is 9.8.

OpenStudy (amistre64):

acc due to g is 9.8 velocity = 9.8t position = 9.8t^2/2 = 4.9 t^2 but since gravity is a downward force; -4.9 t^2

OpenStudy (istim):

Makes sense.

OpenStudy (istim):

With that explained, I think we are done.

OpenStudy (amistre64):

but .... im too old to die!!

OpenStudy (istim):

"You're never to young to die"-A quote that makes me sad.